Suppose the MPC is 0.60. Assume there are no crowding out or investment accelerator effects. If the government increases expenditures by $200 billion, then by how much does aggregate demand shift to the right? If the government decreases taxes by $200 billion, then by how much does aggregate demand shift to the right?

A. $500 billion and $300 billion
B. $500 billion and $500 billion
C. $300 billion and $300 billion
D. $300 billion and $180 billion
Answer #1

Answer: A. $500 billion and $300 billion

Given:

MPC = 0.60 and Government increases expenditures by $200 billion

Calculation:

Spending multiplier = 1 - MPC 1-0.60 = 2.5 0.4

Change in aggregate demand = Initial change in government spending * Spending multiplier

Change in aggregate demand = $200 * 2.5 = $500 billion

Aggregate demand shift to the right by $500 billion

----------------------------------------------------------------------------------------------------------------------------------------------------------------

Given:

MPC = 0.60 and Government decreases taxes by $200 billion

Calculation:

- MPC Tac multiplier == 1- MPC -0.60 1-0.60 -0.60 0.40 = (-)1.5

Change in aggregate demand = Initial change in government tax * Tax multiplier

Change in aggregate demand = $200 * (-) 1.5 = (-) $300 billion

Aggregate demand shift to the right by $300 billion

( Assume there are no crowding out or investment accelerator effects)